### v2

@@ Abstract
Our approach to win first place is divided into two main steps. First, we generate promising initial solutions using two strategies: one based on integer linear programming and the other on employing a combination of geometric greedy heuristics. In the second step, we enhance these solutions through local search techniques, which involve repositioning items and exploring potential replacements to improve the total value of the packing.
@@ 1 Introduction
The CG:SHOP Challenge is an annual competition in geometric optimization. In its sixth edition in 2024, the challenge focuses on a two-dimensional knapsack packing problem. The team Shadoks won first place with the best solution (among the 14 participating teams) to $75$ instances out of $180$ instances.
@@
A total of 180 instances were provided, ranging from $28$ to $50{,}000$ items. The instances are categorized into four classes: Atris, Satris, Jigsaw, and Random, based on the shapes of the items. The Jigsaw instances contain convex polygons with edges in random directions as items and convex containers. The Random instances contain arbitrary polygons as items and convex containers. The Atris instances contain polyominoes as items and rectangular containers. The Satris instances contain polygons obtained by shearing polyominoes as items and rectangular containers. The item values also vary. Some instances have all items with unit value, other instances have values that are small random integers, and other instances have values that are roughly proportional to the area. Solutions to an instance of each class are illustrated in Figure 1, and additional details about the challenge can be found in the organizers’ survey paper [3].
Our general strategy consists of finding good initial solutions (using integer programming or a greedy heuristic) and subsequently optimizing them with local search. Our strategy shares a geometric greedy approach with the second-place team [5], but they did not use integer programming to obtain initial solutions, and their optimization phase is also different from ours. The third-place team [4] uses a hierarchical grid approach. The fourth-place team [1] employed a completely different integer programming model and a genetic algorithm.
We describe the algorithms in Section 2 and experimentally analyze their performance using different parameters in Section 3. Our solvers were coded in Python and C++ and executed on several desktop and laptop computers, as well as the LIMOS and LIS clusters.
@@ 2 Algorithms
We used two different algorithms to compute initial solutions, a preprocessing phase that can be executed beforehand, and a local search phase to improve the solutions.
@@ 2.1 Integer Programming Approach
A simple idea to solve the challenge problem is to produce a set $V$ of random translations of each item inside the container and then reduce the problem to a kind of maximum weight independent set problem in a graph $G=(V,E)$. Each translated item is a vertex with vertex weights determined by the values of the items. There are two types of edges: (1) an edge between two translations that overlap and (2) translations of the same item $i$ form a clique $C_{i}$. If all item have quantity one, then this is a traditional maximum weight independent set problem. However, if items have non-unit quantities, then each clique $C_{i}$ is associated with the quantity $q_{i}$ of item $i$ and at most $q_{i}$ vertices of the clique are allowed in the solution.
This combinatorial problem can easily be modeled as integer programming with one binary variable per vertex. A type-1 edge $uv$ is modeled as $u+v\leq 1$ and each clique $C_{i}$ is modeled as $\sum_{v\in C_{i}}v\leq q_{i}$. The CPLEX solver [2] can optimally solve graphs with a few thousand vertices obtained from the challenge instances, which is not enough to obtain good solutions using uniformly random placements.
To obtain better solutions, we start from a solutions $S$ obtained with the aforementioned method and build a graph $G=(V,E)$ using the same edge rules as before, but for a different set of vertices $V$ constructed as follows. Let $\sigma>0$ be a parameter and $N$ be a set of the zero vector and random vectors where each random vector has $x$ and $y$ coordinates as Gaussian random variables of average 00 and standard deviation $\sigma$. For each item that is placed in $S$ and for each translation vector in $N$, we create a vertex in $V$. Translations that are not completely inside the container are removed. We also create vertices in $V$ using uniform random translations for all items. Edges and cliques are created as before, and the new combinatorial problem is solved with CPLEX. We repeat this procedure multiple times using the previous solution $S$ and reducing the value of $\sigma$ at each step.
@@ 2.2 Greedy Heuristic
The greedy heuristic begins by generating an initial list $L$ of $n$ grid points within the container, where typically $50001000<n<$ 5000. The list $L$ is then shuffled, and its centroid $c$ is computed and rounded to integer coordinates. The point $c$ is subsequently inserted at the beginning of $L$.
Input items are arranged in a list $I$, sorted by decreasing utility. The utility function, detailed in Section 3, is designed to prioritize items with small area and high value. Some example of utility functions we used are the item’s value, the item’s value divided by its area, or the item’s value raised to the power of $1.5$ divided by its area. By starting with high-utility items, the algorithm aims to maximize value while minimizing occupied area. Although this strategy seems intuitive, alternative approaches were explored, such as placing small items at the end of the list to fill remaining spaces as the container becomes full or during local search. However, determining the threshold size for this approach proved challenging, and the idea was ultimately abandoned due to the challenge time constraints.
Regardless of the sorting criteria used, the packing is then constructed by considering items from the list $I$ one by one in order.
At each step, we have a current packing and a new item $i$ to pack. We first try to pack $i$ at the first position $g\in L$, namely in the center of the container. If it overlaps a previously packed item, we try some random positions around $g$. If we do not find a valid position for item around $g$, we move to the next position $g$ in the list $L$ and try some random positions around it. Then the routine that we repeat at each position $g$ of $L$ is the following: first try to pack item $i$ at the grid position $g\in L$. If there are overlaps with the previously packed items or the boundary of the container, try some random positions around $g$. If no valid positions is found for item $i$ around $g$, skip to the next position.
@@
We select a direction $u$ along which the item $i$ is pushed until it can no longer advance due to obstructions from either other packed items or the boundaries of the container. Pushing an item entails translating it in directions $v$ such that the dot product $v.u$ with $u$ is strictly positive. Since the dot product of the item’s position with $u$ always increases during the process, the possibility of loops is avoided. The push routine terminates when no further movement is possible in any direction $v$.
There are several strategies for selecting the push direction $u$ (see Section 3 for details). The most efficient approach is to push the items in a direction normal to their diameter. Alternatively, a separation criterion can be introduced, where thinner items are pushed to the left and rounder items to the right. The greedy algorithm ends when all items in $I$ have either been packed or tested at all positions in the list $L$. Four distinct strategies are illustrated in Figure 2.
@@ 2.3 Clusters Preprocessing
Our greedy heuristic does not mimic the human approach to solve small packing instances. When faced with a packing puzzle, a human would likely first group items into compact clusters before attempting to pack them into the container. In other words, a human would preprocess the items to identify useful clusters.
The objective of the preprocessing step is to combine small sets of items into larger groups. The term cluster refers to a set of items with fixed relative positions. A cluster functions like an individual item but is a polygonal set rather than a simple polygon.
@@
cluster utility($\text{cluster utility}(cluster)=\frac{gauss\cdot value^{\alpha}\cdot penalty}{area}$
@@
  • $gauss$ is a random Gaussian variable with mean $1$ to diversify clusters upon repeated computation,
  • $value$ is the sum of item values,
@@
  • $area$ is either the convex hull area or a weighted sum with items areas,
  • $penalty$ is fixed equal to $1$ except in some cases presenting clusters that we would like to discourage. Let us consider for instance that we have $10$ copies of an axis-parallel rectangular of width $10$ and height $1$. There are two ways to assemble the copies of this item: either in a long bar of size $100\times 1$ or in a square of size $10\times 10$. The long bar does not seem a good choice but without penalty, its cluster utility would be equal to the one of the square. Then we penalize thin clusters from a factor $penalty<1$ computed by using the thickness of the items and the thickness of the cluster.
@@
With thousands of items, evaluating all possible pairs for assembly is infeasible. Thus, the first task is to construct a compatibility graph connecting items that can potentially be assembled together. But instead of computing a unique compatibility graph, we compute several according different criterion. It provides a better control on the clusters generation.
The second step consists in generating clusters from a compatibility graph. We start by generating clusters with $2$ items until a fixed maximum number of items. The items are the generation $1$ and the generation $k+1$ (clusters with $k+1$ items) is built from the generations $1$ and $k$. The clusters of the generation $k+1$ items are computed by assembling clusters of the generation $k$ with items which are compatible to at least one of their own items. This implies that the items within a cluster of generation $k$ form a connected component of size $k$ in the compatibility graphs. If the average degree of the compatibility graphs is larger than a small constant, it becomes impractical to attempt all possibilities. A crucial requirement is to manage the exponential growth of the number of potential clusters. We achieve this goal in all generations by keeping only a fixed number $m$ of clusters per item. The selection is done according to the cluster utility function. For each item, we keep the $m$ clusters of highest cluster utility.
@@
  • A second compatibility graph is denoted Thin. This graph connects the items whose ratio diameter/width is larger than a constant (e.g., $3$) and whose longest edges have roughly the same direction.
@@
  • Satris items are sheared polyominoes. We build a compatibility graph called Shear by taking account the directions of the edges of the items. Each item has two dominant directions. We group the items with close dominant directions and take as Shear the union of the cliques of these groups.
  • Atris items are slightly pertubed polyominoes but the perturbations are sufficiently small to allow us to encode each shape by its Freeman code. Then the complementarity of item pairs can easily checked by word combinatoric. Nevertheless, the instances of Atris used only a small number of shapes (bars, crosses, L, shapes in V and waves). It makes possible to achieve a complete classification of the items and thus build a compatibility graph which connects all the items of a given class to all the items of another one. This compatibility graph is denoted Polyo.
@@
  • Vertex-based approach: Consider all pairs of item and cluster vertices. Translate the item vertex onto the cluster vertex. If there are no overlaps, then compute the new cluster utility.
@@
Unsurprisingly, clusters work particularly well for the atris instances, where items resemble polyominoes. During the CG 2024 competition, the computation time for clusters was not sufficiently controlled, limiting the use of this strategy for large instances. Consequently, we only made moderate use of clusters.
@@ 2.4 Local Search
The strategy of the dig routine is simply to choose a point $v$ of the container and to push the packed items as far as possible from this point in order to free some space. Then the dig routine first randomly chooses either a point $v$ in the interior of the container or a vertex of the container. The packed items are sorted by the distance of their centroids to the point $v$, from the farthest to the closest. Then, each packed item $i$ with centroid $c_{i}$ is pushed in the direction $c_{i}-v$. The goal is to free the space around $v$. Once all packed items have been pushed, we try to pack the unpacked items around $v$ by using the grid points $g$ in $L$ close to $v$ and some random integer points around each point $g$. It may be the case that an item can be packed if we remove some other items of the packing that intersect it. In this case, we compute the benefit of removing the conflicting items in order to pack the new item, replacing the item if the benefit is non-negative. To accelerate the procedure for large instances, we add a parameter which allows us to push only the items that are close to $v$.
@@ 3 Parameter Analysis
In order to compare the different parameters, we consider $18$ instances of various sizes, as shown in Table 1. Throughout, we refer to the ratio between the value of the solution in question and the value of the best solution in the challenge as the value ratio (notice that the competition score is the square of the value ratio). The best solutions we found to the first $15$ table instances are shown in Figure 4.
Instance IP IP + LS Gr Gr $+$ LS Our best
jigsaw_cf4_273db689_28 $\mathbf{1}$ $\mathbf{1}$$900$950.950.950.950 $1$
random_cf1_64ac4991_50 .$\mathbf{.926}$$\mathbf{.926}$$851$$925$926.926.926.926
jigsaw_rcf4_7702a097_70 .$959$$\mathbf{.982}$$867$924.924.924.924 $1$
random_cf4_50e0d4d9_100 .$943$$\mathbf{.972}$$877$941.941.941.941 $1$
random_cf3_x21f5def_200 .$950$$\mathbf{.967}$$893$$\mathbf{.967}$ $1$
random_rcf1_340f4443_500 .$952$$960$$932$$\mathbf{.981}$984.984.984.984
random_rcf3_x7651267_1000 .$927$$970$$926$$\mathbf{.980}$ $1$
jigsaw_rcf4_6de1b3b7_1363 .$934$$948$$928$$\mathbf{.980}$ $1$
random_cf1_x51ab828_2000 .$937$$945$$892$$\mathbf{.950}$961.961.961.961
atris2986 .$910$$931$$881$$\mathbf{.948}$988.988.988.988
atris3323 .$893$$918$$866$$\mathbf{.951}$ $1$
satris4681 .$919$$930$$886$$\mathbf{.937}$ $1$
jigsaw_cf1_4fd4c46e_6548 .944.944.944.944 .944.944.944.944 .$955$$\mathbf{.963}$975.975.975.975
atris7260 .$886$$902$$872$$\mathbf{.920}$974.974.974.974
random_cf3_x4b49fe2_10000 .$924$$938$$889$$\mathbf{.963}$ $1$
satris15666 .$923$$\mathbf{.951}$$877$910.910.910.910 $1$
jigsaw_cf1_$2$$732$$876$$972$$\mathbf{.973}$985.985.985.985
atris41643 .$513$$803$$854$$\mathbf{.877}$911.911.911.911
Table 1: Value ratio of several instances using integer programming (IP) or greedy (Gr), both before and after 24 hours of local search (LS).
@@
Figure 5: The value ratio over time for the instances random_cf4_50e0d4d9_100, random_cf3_x21f5def_200, and random_rcf3_x7651267_1000 for $12$ independent executions of the greedy algorithm and a subsequent local search. Red curves are computed without cluster preprocessing, while the light blue and dark blue curves respectively include a preprocessing of clusters with at most $5$ and $10$ items.
@@
  • The number of grid positions and random integer positions considered around each point varies. Typically, we used between $500$ and $3000$ grid positions, along with $3$ to $10$ random positions per grid point. Figure 6 compares different parameter settings and shows that using more than $1000$ grid positions offers no significant improvement.
@@
  • The direction $u$ in which an item is pushed can be selected in different ways. It may be chosen randomly (strategy 1) or as a normal to the item’s diameter, with a random choice of which side to push (strategy 2). However, the following method was primarily used: if the item is thin (with a diameter-to-width ratio greater than $3$), it is pushed to the left along the normal to its diameter; if the item is fat (ratio less than $3$), it is pushed to the right (strategy 3). Additionally, items may be pushed in the direction normal to their longest edge if they are fat (strategy 4). Figure 8 shows the average results over 10 runs for each instance from Table 1.
@@
$value$               $value/area$           $value^{1.5}/area$        $(1+t)value/area$
@@
Figure 8: Value ratio of the greedy algorithm (orange) using different push strategies, for the 18 instances in Table 1. In red, we show the value attained after 1 hour of local search.
Choosing the best parameters still remains a challenging task.

### v3

@@ Abstract
Our approach to win first place is divided into two main steps. First, we generate promising initial solutions using two strategies: one based on integer linear programming and the other one employing a combination of geometric greedy heuristics. In the second step, we enhance these solutions through local search techniques, which involve repositioning items and exploring potential replacements to improve the total value of the packing.
@@ 1 Introduction
The CG:SHOP Challenge is an annual competition in geometric optimization. In its sixth edition in 2024, the challenge focuses on a two-dimensional knapsack packing problem. Our team, called Shadoks, won first place with the best solution (among the 14 participating teams) to $75$ instances out of $180$ instances.
@@
A total of 180 instances were provided, ranging from $28$ to 50,00050{,}000 items. The instances are categorized into four classes: Atris, Satris, Jigsaw, and Random, based on the shapes of the items. The Jigsaw instances contain convex polygons with edges in random directions as items and convex containers. The Random instances contain arbitrary polygons as items and convex containers. The Atris instances contain slightly perturbed polyominoes as items and rectangular containers. The Satris instances contain polygons obtained by shearing polyominoes as items and rectangular containers. The item values also vary. Some instances have all items with unit value, other instances have values that are small random integers, and other instances have values that are roughly proportional to the area. Solutions to an instance of each class are illustrated in Figure 1, and additional details about the challenge can be found in the organizers’ survey paper [survey].
Our general strategy consists of finding good initial solutions (using integer programming or a greedy heuristic) and subsequently optimizing them with local search. Our strategy shares a geometric greedy approach with the second-place team [place2], but they did not use integer programming to obtain initial solutions, and their optimization phase is also different from ours. The third-place team [place3] uses a hierarchical grid approach. The fourth-place team [place4] employed a completely different integer programming model and a genetic algorithm.
We describe our algorithms in Section 2 and experimentally analyze their performance using different parameters in Section 3. Section 4 presents engineering details focusing on performance. Our solvers were coded in Python and C++ and executed on several desktop and laptop computers, as well as the clusters available in our labs (LIS and LIMOS).
@@ 2 Algorithms
We use two different algorithms to compute initial solutions, a preprocessing phase that can be executed beforehand, and a local search phase to improve the solutions.
@@ 2.1 Integer Programming Approach
A simple idea to solve the challenge problem is to produce a set $V$ of random translations of each item inside the container and then reduce the problem to a type of maximum weight independent set problem in a graph $G=(V,E)$. The vertex set $V$ is a set of translated items. The weight associated to each vertex $v$ is the value of the corresponding item. The edge set $E$ contains two types of edges. Two vertices are connected by an edge of the first type if their translated items overlap. Hence, an independent set will not contain overlapping items. The second type of edge enforces that the solution does not have more than the prescribed quantity $q_{i}$ of a given item $i$. To take this constraint into account, we introduce a clique $C_{i}$ connecting all the vertices corresponding to item $i$. If the quantity $q_{i}$ is $1$, then we need to choose at most one vertex per clique. If all quantities are $1$, then the packing problem (restricted to a finite set of translations) is equivalent to the traditional maximum weight independent set problem. However, if items have non-unit quantities, then each clique $C_{i}$ is associated with the quantity $q_{i}$ of item $i$ and at most $q_{i}$ vertices of the clique are allowed in the solution.
The combinatorial problem described in the previous paragraph is easily modeled as integer programming with one binary variable per vertex. A type-1 edge $uv$ is modeled as $x_{u}+x_{v}\leq 1$ where $x_{u}$ and $x_{v}$ are, respectively, the variables associated to the vertices $u$ and $v$. Each clique $C_{i}$ is modeled as $\sum_{v\in C_{i}}x_{v}\leq q_{i}$. The CPLEX solver [cplex] can optimally solve graphs with a few thousand vertices obtained from the challenge instances, which is not enough to obtain good solutions using uniformly random translations.
To obtain better solutions, we start from a solution $S$ obtained with the aforementioned method and build a graph $G=(V,E)$ using the same edge rules as before, but for a different set of vertices $V$ constructed as follows. Let $\sigma>0$ be a parameter and $N$ be a set of the zero vector and random vectors where each random vector has $x$ and $y$ coordinates as Gaussian random variables of mean 0 and standard deviation $\sigma$. For each translated item in the solution $S$ and for each translation vector in $N$, we create a vertex in $V$ by translating the item accordingly. Translations that are not completely inside the container are removed. We also create vertices in $V$ using uniformly random translations for all items. Edges and cliques are created as before, and the new combinatorial problem is solved with CPLEX. We repeat this procedure multiple times using the previous solution $S$ and reducing the value of $\sigma$ at each step.
@@ 2.2 Greedy Heuristic
The greedy heuristic begins by generating an initial list $L$ of $n$ grid points within the container, where typically $1000<n<5000$. List $L$ is then shuffled, and its centroid $c$ is computed and rounded to integer coordinates. The point $c$ is subsequently inserted at the beginning of $L$.
Input items are arranged in a list $I$, sorted by decreasing utility. The utility function, detailed in Section 3, is designed to prioritize items with small area and high value. Some example of utility functions we use are the item’s value, the item’s value divided by its area, or the item’s value raised to the power of $1.5$ divided by its area. By starting with high-utility items, the algorithm aims to maximize value while minimizing occupied area. Although this strategy seems intuitive, alternative approaches were explored, such as placing small items at the end of the list to fill remaining spaces as the container becomes full or during local search. However, determining the threshold size for this approach proved challenging, and the idea was ultimately abandoned due to the challenge time constraints.
Regardless of the sorting criteria used, the packing is then constructed by considering items from list $I$ one by one in order.
At each step, we have a current packing and a new item $i$ to pack. We first try to pack $i$ at the first position $g\in L$, namely in the center of the container. If it overlaps a previously packed item, we try some random positions around $g$. If we do not find a valid position for item $i$ around $g$, we move to the next position $g$ in list $L$ and try some random positions around it. Then the routine that we repeat at each position $g$ of $L$ is the following: first try to pack item $i$ at grid position $g\in L$. If there are overlaps with the previously packed items or the boundary of the container, try some random positions around $g$. If no valid position is found for item $i$ around $g$, skip to the next position.
@@
We select a direction $u$ along which item $i$ is pushed until it can no longer advance due to obstructions from either other packed items or the boundary of the container. Pushing an item entails translating it in directions $v$ such that the dot product $v.u$ with $u$ is strictly positive. Since the dot product of the item’s position with $u$ always increases during the process, the possibility of loops is avoided. The push routine terminates when no further movement is possible in any direction $v$.
There are several strategies for selecting the push direction $u$ (see Section 3 for details). The most efficient approach is to push the items in a direction normal to their diameter. Alternatively, a separation criterion can be introduced, where skinny items are pushed to the left and fat items to the right. The greedy algorithm ends when all items in $I$ have either been packed or tested at all positions in list $L$. Four distinct strategies are illustrated in Figure 2.
@@ 2.3 Clusters Preprocessing
Our greedy heuristic is a first approach which mimics the human approach to solve packing instances but there is a second one. When faced with a packing puzzle, a human would likely first group items into compact clusters before attempting to pack them into the container. In other words, a human would preprocess the items to identify useful clusters.
The objective of the preprocessing step is to combine small sets of items into larger groups. The term cluster refers to a set of items with fixed relative positions. A cluster functions like an individual item but is a polygonal set rather than a simple polygon (Fig. 3).
Refer to captionRefer to caption
@@
$\text{cluster utility}(cluster)=\frac{gauss\cdot value^{\alpha}\cdot penalty}{area}$
@@
  • gaussgauss is a random Gaussian variable with mean $1$ to diversify clusters upon repeated computation,
  • valuevalue is the sum of item values,
@@
  • areaarea is either the convex hull area or a weighted sum with item areas,
  • penaltypenalty is fixed equal to $1$ except for some cases presenting clusters that we would like to discourage. Let us consider for instance that we have $10$ copies of an axis-parallel rectangle of width $10$ and height $1$. There are two ways to assemble the copies of this item: either in a long bar of size $100\times 1$ or in a square of size $10\times 10$. The long bar does not seem a good choice but without penalty, its cluster utility would be equal to the one of the square. Then we penalize skinny clusters from a factor $penalty<1$ computed by using the thickness of the items and the thickness of the cluster.
@@
With thousands of items, evaluating all possible pairs for assembly is infeasible. Thus, the first task is to construct a compatibility graph connecting items that can potentially be assembled together. But instead of computing a single compatibility graph, we compute multiple graphs, each based on different criteria. It provides a better control on the cluster generation.
The second step consists in generating clusters from a compatibility graph. We start by generating clusters with two items until a fixed maximum number of items. The items are generation $1$ and generation $k+1$ (clusters with $k+1$ items) is built from generations $1$ and $k$. The clusters of generation $k+1$ items are computed by assembling clusters of generation $k$ with items which are compatible to at least one of their own items. This implies that the items within a cluster of generation $k$ form a connected component of size $k$ in the compatibility graphs. If the average degree of the compatibility graphs is larger than a small constant, it becomes impractical to attempt all possibilities. A crucial requirement is to manage the exponential growth of the number of potential clusters. We achieve this goal in all generations by keeping only a fixed number $m$ of clusters per item. The selection is done according to the cluster utility function. For each item, we keep the $m$ clusters of highest cluster utility.
@@
  • A second compatibility graph is denoted Skinny. This graph connects the items whose ratio diameter/width is larger than a constant (e.g., $3$). It connects the items whose longest edges have roughly the same direction.
@@
  • Satris items are sheared polyominoes. We build a compatibility graph called Shear by taking the directions of the edges of the items into account. Each item has two dominant directions. We group the items with close dominant directions and take as Shear the union of the cliques of these groups.
  • Atris items are slightly perturbed polyominoes but the perturbations are sufficiently small to allow us to encode each shape by its contour path i.e a word on the alphabet $\{(1,0),(0,1),(-1,0),(0,-1)\}$. Then the complementarity of item pairs can easily be checked by word combinatorics. Nevertheless, the instances of Atris use only a small number of shapes (bars, crosses, L, shapes in Y and waves). The classification of the items allows us to build a compatibility graph that links all elements of a given class to all the elements of a class with a complementary shape.
@@
  • Vertex-based approach: Consider all pairs of an item vertex and a cluster vertex. Adjust the relative positions of the cluster and the item so that the item vertex is translated onto the cluster vertex. If there are no overlaps, then compute the new cluster utility.
@@
Unsurprisingly, clusters work particularly well for the atris instances, where items resemble polyominoes. During the CG 2024 competition, our cluster-based approach faced scalability limitations: computation time was not adequately controlled, rendering the strategy insufficiently robust for large instances. As a result, clusters were only sparingly employed in our overall solution, highlighting the need for further work on the subject.
@@ 3 Parameter Analysis
In order to compare the different parameters, we consider $18$ instances of various sizes, as shown in Table 1. Throughout, we refer to the ratio between the value of the solution and the value of the best solution in the challenge as the value ratio (notice that the competition score is the square of the value ratio). The best solutions we found to the first $15$ table instances are shown in Figure 4.
Instance IP IP + LS Gr Gr ++ LS Our best
jigsaw_cf4_273db689_28 $\mathbf{1}$ $\mathbf{1}$ .900.900 .950.950 11
random_cf1_64ac4991_50 $\mathbf{.926}$ $\mathbf{.926}$ .851.851 .925.925 .926.926
jigsaw_rcf4_7702a097_70 .959.959 $\mathbf{.982}$ .867.867 .924.924 11
random_cf4_50e0d4d9_100 .943.943 $\mathbf{.972}$ .877.877 .941.941 11
random_cf3_x21f5def_200 .950.950 $\mathbf{.967}$ .893.893 $\mathbf{.967}$ 11
random_rcf1_340f4443_500 .952.952 .960.960 .932.932 $\mathbf{.981}$ .984.984
random_rcf3_x7651267_1000 .927.927 .970.970 .926.926 $\mathbf{.980}$ 11
jigsaw_rcf4_6de1b3b7_1363 .934.934 .948.948 .928.928 $\mathbf{.980}$ 11
random_cf1_x51ab828_2000 .937.937 .945.945 .892.892 $\mathbf{.950}$ .961.961
atris2986 .910.910 .931.931 .881.881 $\mathbf{.948}$ .988.988
atris3323 .893.893 .918.918 .866.866 $\mathbf{.951}$ 11
satris4681 .919.919 .930.930 .886.886 $\mathbf{.937}$ 11
jigsaw_cf1_4fd4c46e_6548 .944.944 .944.944 .955.955 $\mathbf{.963}$ .975.975
atris7260 .886.886 .902.902 .872.872 $\mathbf{.920}$ .974.974
random_cf3_x4b49fe2_10000 .924.924 .938.938 .889.889 $\mathbf{.963}$ 11
satris15666 .923.923 $\mathbf{.951}$ .877.877 .910.910 11
jigsaw_cf1_203072aa_32622 .732.732 .876.876 .972.972 $\mathbf{.973}$ .985.985
atris41643 .513.513 .803.803 .854.854 $\mathbf{.877}$ .911.911
Table 1: Value ratio of several instances using integer programming (IP) or greedy (Gr), both before and after 24 hours of local search (LS). The last column shows the best value ratio that the Shadoks team obtained in the competition.
@@
Figure 5: The value ratio over time for the instances random_cf4_50e0d4d9_100, random_cf3_x21f5def_200, and random_rcf3_x7651267_1000 for $12$ independent executions of the greedy algorithm and a subsequent local search. Red curves are computed without cluster preprocessing, while the light blue and dark blue curves respectively include a preprocessing of clusters with at most 55 and $10$ items.
@@
  • The number of grid positions and random integer positions considered around each point varies. Typically, we use between $500$ and $3000$ grid positions, along with $3$ to $10$ random positions per grid point. Figure 6 compares different parameter settings and shows that using more than $1000$ grid positions offers no significant improvement.
@@
  • The direction $u$ in which an item is pushed can be selected in different ways. It may be chosen randomly (strategy 1) or as a normal to the item’s diameter, with a random choice of which side to push (strategy 2). If the item is skinny (with a diameter-to-width ratio greater than $3$), then it is pushed to the left along the normal to its diameter; if the item is fat (ratio less than $3$), then it is pushed to the right (strategy 3). Additionally, items may be pushed in the direction normal to their longest edge if they are fat (strategy 4). Figure 8 shows the average results over 10 runs for each instance from Table 1. According to our experiments, the best strategy is strategy 3. Additional experiments could have been done by using directions other than just left and right as for instance up and down but we assumed that the results would be similar for the instances of the challenge due to the shapes of the containers.
@@
     valuevalue        $value/area$      $value^{1.5}/area$   $(1+t)value/area$
@@
Figure 8: Value ratio of the greedy algorithm (orange) using different push strategies for the 18 instances in Table 1. In red we show the value attained after 1 hour of local search.
Since our experiments to identify the optimal utility function yielded inconclusive results, a pragmatic strategy is to reduce the grid resolution to accelerate the greedy algorithm’s execution in order to broader test of alternative utility functions. Nevertheless, selecting the best parameters for a given instance remains a non-trivial problem.